# Parametrized Sharing for Multi-Agent Hybrid DRL for Multiple Multi-Functional RISs-Aided Downlink NOMA Networks

### v1

@@ II System Model and Problem Formulation
$\mathbf{w}_{q}\in\mathcal{W},\quad\qquad\qquad\forall q\in\mathcal{Q},$ (4g)
where $P_{\text{total}}=\sum_{q\in\mathcal{Q}}(P_{q}^{\text{con}}-\sum_{m\in\mathcal{M}}P_{q,m}^{\text{A}})+\sum_{k\in\mathcal{K}}\left\lVert\mathbf{f}_{k}\right\rVert^{2}$ is the system total consumed power. Constraint set in $\mathcal{R}_{\mathbf{\Theta}}$ (4b) specifies the feasible region of MF-RISs, i.e., $\alpha_{q,m}\in[0,1]$, $\beta_{q,m}^{k}\in[0,\beta_{\max}^{k}]$, $\theta_{q,m}^{k}\in[0,2\pi)$. Constraint (4c) ensures the minimum rate requirement per user as $R^{\min}_{kj}$ while constraint (4d) represents the NOMA power allocation restriction. Constraint (4e) ensures that the total BS transmit power cannot exceed its budget $P_{BS}^{\max}$. Due to non-convexity and non-linearity of problem (4a), it presents a significant challenge to solve this problem. To address these difficulties, we propose a DRL-based scheme, which is detailed in the following section.
@@ III-A Hybrid DRL Algorithm
  • Reward: We design the shared reward as the overall EE in conjunction with its constraints as penalties, given by
    $r(t)=\frac{\sum_{k\in\mathcal{K}}\sum_{j\in\mathcal{J}_{k}}R_{kj}}{P_{\text{total}}}-\sum_{i=1}^{3}\rho_{i}C_{i},$ (5)
    where $\rho_{i},\forall i\in\{1,2,3\}$ indicates the weights of each penalty $C_{i}$ corresponding to constraints of (4c), (4e), and (4f), which are defined as $C_{1}=\sum_{k\in\mathcal{K}}\sum_{j\in\mathcal{J}_{k}}(R_{kj}^{\min}-R_{kj})$, $C_{2}=\sum_{k\in\mathcal{K}}\mathbf{\|}\mathbf{f}_{k}\|^{2}-P_{BS}^{\max}$, and $C_{3}=\sum_{q\in\mathcal{Q}}(P_{q}^{\text{con}}-\sum_{m\in\mathcal{M}}P_{q,m}^{\text{A}})$, respectively. Note that the remaining boundary conditions in (4b), (4d), and (4g) can be automatically constrained during generating actions.
@@ III-A1 DQN for Discrete Variables
DQN employs a deep neural network, Q-network, to approximate the Q-function $Q_{q}(s_{q},a_{q}|\omega_{q}^{\phi})$ which estimates the expected cumulative reward for each action $a_{q}$ under a given state $s_{q}$. Note that we define $\omega^{\phi}_{q}$ and $\omega^{\phi^{-}}_{q}$ as the model weights of the current network and of the target network of DQN, respectively. Based on estimated Q-values, each agent selects its action using an $\epsilon$-greedy strategy, which balances exploration and exploitation by choosing a random action with probability $\epsilon$ and by selecting the action with the maximum predicted Q-value with probability $1-\epsilon$, i.e., $\epsilon(t)=\epsilon({t-1})-\frac{\epsilon_{\max}-\epsilon_{\min}}{\epsilon_{d}}$, where $\epsilon_{d}$ is the decay parameter. Notations of $\epsilon_{\max}$ and $\epsilon_{\min}$ indicate the maximum and minimum exploration boundaries, respectively. To enhance training stability, DQN incorporates two critical techniques: (i) Experience replay buffer stores historical trajectories with a tuple of $(s_{q},a_{q},r_{q},s^{\prime}_{q})$, allowing the agent to sample mini-batches uniformly and eliminate temporal correlations during learning, where $s^{\prime}_{q}$ indicates the new state; and (ii) Target network, with its model denoted as $Q^{\prime}_{q}(s_{q},a_{q}|\omega_{q}^{\phi^{-}})$ is periodically softly updated to provide stable Q-learning, i.e., $\omega_{q}^{\phi^{-}}\leftarrow\tau_{\phi}\omega_{q}^{\phi}+(1-\tau_{\phi})\omega_{q}^{\phi^{-}}$ where $\tau_{q}^{\phi}$ indicates the importance of target model of DQN. The Q-network is then updated by minimizing the temporal-difference (TD) error, which measures the discrepancy between the predicted Q-value and the target Q-value, given by
$\mathcal{L}(\omega_{q}^{\phi})=\mathbb{E}_{(s_{q},a_{q},r,s^{\prime}_{q})}[(y-Q_{q}(s_{q},a_{q}\mid\omega_{q}^{\phi}))^{2}],$ (6)
where $y=r(t)+\gamma^{d}\max_{a^{\prime}}Q^{\prime}_{q}(s^{\prime}_{q},a^{\prime}\mid\omega_{q}^{\phi^{-}})$ indicates the TD target value and the discount factor $\gamma^{d}\in[0,1]$ indicates the importance of future rewards. Notation of $Q^{\prime}_{q}(\cdot)$ is the Q-value of the target network. The parameter update via gradient descent is then given by $\omega_{q}^{\phi}\leftarrow\omega_{q}^{\phi}-l^{d}\cdot\nabla_{\omega_{q}^{\phi}}\,\mathcal{L}(\omega_{q}^{\phi})$, where $l^{d}\in[0,1]$ is the learning rate.
@@ III-A2 PPO for Continuous Variables
The remaining continuous parameters are optimized using the PPO algorithm [ref21]. Particularly, PPO adopts an actor-critic framework respectively consisting of a policy network and of a value network. In the policy network, the neural network outputs the mean and standard deviation of a multivariate Gaussian distribution, from which actions are sampled according to the current state $s_{q}(t)$ and policy $\pi_{\delta_{q}}(a_{q}(t)|s_{q}(t))$. Note that $\delta_{q}$ indicates the policy network parameters. To optimize the policy network, we employ a clipped surrogate objective function expressed as
$\mathcal{L}^{\text{clip}}(\delta_{q})={\mathbb{E}}_{t}$ $\big[\min\big(O_{q}(\delta_{q})\hat{A}_{q}(t),$
$\quad\mathrm{clip}(O_{q}(\delta_{q}),1-\Lambda,1+\Lambda)\hat{A}_{q}(t)\big)\big],$ (7)
where $\mathbb{E}$$\cdot$] is the expectation over a batch of generated trajectories, and $O_{q}(\delta_{q})=\frac{\pi_{\delta_{q}}(a_{q}(t)|s_{q}(t))}{\pi_{\delta_{q,\text{old}}}(a_{q}(t)|s_{q}(t))}$ is the probability ratio. $\mathrm{clip}(\cdot)$ indicates the clipping function which clips the change between the new and old policies within the range $1-\Lambda,1+\Lambda]$ for avoiding excessive policy updates. Note that $\delta_{q,\text{old}}(\cdot)$ is the old policy parameters. Furthermore, $\hat{A}_{q}(t)$ is the generalized advantage estimation (GAE) quantifying the difference between the observed outcome of each action in a state and the predicted state value $V_{\mu_{q}}(s_{q}(t))$ by the value network, which is $\hat{A}_{q}(t)\!=\!\sum_{i=t}^{T-t}(\gamma^{p}\lambda^{p})^{i-t}\{[(r(i)\!+\!\gamma^{p}V_{\mu_{q}}(s_{q}(i\!+\!1))]\!-\!V_{\mu_{q}}(s_{q}(i))\}$, where $\gamma^{p}$ and $\lambda^{p}$ are importance ratio and GAE hyperparameters, respectively. Notation $T$ means the length of trajectory segment. The policy is optimized iteratively by maximizing the clipped surrogate objective using gradient ascent given by $\delta_{q}\leftarrow{\delta_{q}}-l^{p}_{\text{ac}}\cdot\nabla_{\delta_{q}}\mathcal{L}^{\text{clip}}(\delta_{q})$, where $l^{p}_{\text{ac}}\in[0,1]$ is learning rate for actor in PPO. The associated loss function of the value network is defined as $\mathcal{L}^{V}(\mu_{q})=\mathbb{E}_{t}[\big(V_{\mu_{q}}(s_{q}(t))-\hat{V}^{\text{tar}}_{q}(t)\big)^{2}]$, where $\hat{V}^{\text{tar}}_{q}(t)=V_{\mu_{q}^{-}}\!\big(s_{q}(t)\big)+\hat{A}_{q}(t)$ and $\mu_{q}^{-}$ indicates the previous update of value network. The parameters of value network are updated by the gradient method, i.e., $\mu_{q}\leftarrow{\mu_{q}}-l^{p}_{\text{cr}}\cdot\nabla_{\mu_{q}}\mathcal{L}^{V}(\mu_{q})$, where $l^{p}_{\text{cr}}\in[0,1]$ is learning rate for critic in PPO.
@@ III-B Parametrized Sharing in PMHRL
In the context of individual agent design in hybrid DRL frameworks, PPO and DQN typically select actions independently based on their respective input states. This isolated decision-making process neglects the potential interdependence and interaction between the two strategies. To address this, inspired by [ref23], we propose a parametrized sharing mechanism. The core idea is to enable the shared representation between the PPO and DQN models by exchanging features. Since PPO handling high-dimensional actions performs a more complex task than DQN, information from DQN is essential to PPO. Specifically, PPO agent utilizes the discrete action output from the DQN agent as the input of PPO, i.e., $s_{q}^{\text{con}}(t)=\operatorname{concat}(\mathbf{g}_{q,kj}(t),\operatorname{vec}(a_{q}^{\text{dis}}(t-1)))$, where ${\rm concat}(\cdot)$ indicates the concatenation of two vectors and ${\rm vec}(\cdot)$ vectorizes the discrete action. Note that only MF-RISs require parametrized sharing as they have hybrid actions, thereby improving coordination between the two decision modules.
@@ IV Simulation Results
MF-RIS power consumption parameters [ref19, ref17-1] $\xi=1.1$, $P_{\text{PIN}}=0.33$ mW, $P_{C}=2.1$ mW, $Z=24$ mW, $\varpi_{1}=150$, $\varpi_{2}=0.014$, $L_{\alpha}=2$, $L_{\beta}=10$, $L_{\theta}=8$
@@
Figure 2: (a) Convergence. (b) $E$ with different strategies (c) $E$ with different MF-RIS cases. (d) Comparison between NOMA, SDMA, and OMA.
In simulations, we evaluate PMHRL in multi-MF-RISs-assisted downlink NOMA. We consider the BS positioned at $\mathbf{w}_{b}=[0,0,5]$ m, serving $J_{k}=2$ users in $K=2$ directions. Then the users are randomly distributed within a circular area of radius 2 m, centered at $[0,30,0]$, $[0,35,0]$, $[10,40,0]$, and $[10,45,0]$ m, respectively. The MF-RISs are equipped with $M=32$ elements, and the BS has $N=6$ antennas. The remaining parameters related to networks are listed in Table I. As for PMHRL, learning rates of PPO actor/critic networks are $l^{p}_{\text{ac}}=10^{-3}$ and $l^{p}_{\text{cr}}=10^{-4}$, respectively, whereas that of DQN is $l^{d}=10^{-3}$. The discount factor for both modules is $\gamma^{d}=\gamma^{p}=0.99$. The decay and soft update parameters of DQN are set to $\epsilon_{d}=10^{4}$ and $\tau_{\phi}=10^{-2}$, respectively. The experience replay buffer of DQN stores up to $10^{6}$ samples. The mini-batch mechanism is adopted during training, with a batch size of $64$. $\epsilon_{\max}=1$ and $\epsilon_{\min}=0$. Moreover, we set the clipping ratio to $\Lambda=0.2$, the GAE parameter is $\lambda^{p}=0.97$, and trajectory length is $10^{3}$. The weights of each penalty in (5) are $\rho_{1}=10^{-3}$, $\rho_{2}=\rho_{3}=10^{-5}$.
@@
In Fig. 2, the results show that EE escalates with the increasing numbers of antennas thanks to improved beamforming capability, reaching a peak at $N=6$ before declining as the power consumption begins to outweigh the beamforming gains. In addition, we compare the fully-optimized case to the cases without either EH ratio $\alpha_{q,m}$, amplification $\beta_{q,m}^{k}$, phase-shift $\theta^{k}_{q,m}$, or deployment $\mathbf{w}_{q}$. Note that ”w/o” indicates the random decision. It is evident that the full optimization yields the highest EE. In contrast, omitting the optimization of specific parameters leads to noticeable EE degradation. The configuration without optimizing EH ratio results in the lowest EE. This is because random S or H mode selection leads to inefficient EH, where the collected energy fails to compensate for high power consumption.
Fig. 2 compares EE of $Q\in\{2,3\}$ MF-RISs under different cases: (1) Optimized case; (2) No EH ($\alpha_{q,m}^{k}=\alpha_{q,m}=1,\forall k\in\mathcal{K}$); (3) No amplification ($\beta_{q,m}^{k}=\beta_{q,m}=1,\forall k\in\mathcal{K}$); (4) Only reflection capability. The results show that increasing the number of MF-RIS elements initially enhances the EE. However, further increasing elements induces higher energy consumption, leading to a degraded EE owing to insufficient support from harvested energy. Notably, the case with $Q=3$ MF-RISs outperforms that with $Q=2$ MF-RISs across all cases, attributed to the enhanced spatial diversity and EH gain from multiple MF-RISs. Specifically, when the signal amplification is disabled, the reduced signal gain leads to a lower EE than that of the fully-optimized case. Moreover, in the case of MF-RIS only with reflection, the signal cannot be delivered to users beyond the other side of the surface, leading to a significant EE reduction. Additionally, when the EH function is fully disabled, the system cannot support the extra power from MF-RISs, resulting in the lowest EE among all cases.
Fig. 2 reveals the EE performance under varying numbers of users. We compare NOMA to the existing multiple access mechanisms, i.e., OMA and spatial division multiple access (SDMA) [ref22] with or without deployment of MF-RISs. As observed, EE decreases with more users due to insufficient power resources and severe inter-user interference. Moreover, NOMA with MF-RIS deployment achieves the highest EE across all numbers of users, benefiting from its superior spectrum utilization and EH capabilities offered by MF-RISs to SDMA and OMA. In contrast, their counterparts without deploying MF-RIS show significantly lower EE performance.

### v2

@@ II System Model and Problem Formulation
where $P_{\text{total}}=\sum_{q\in\mathcal{Q}}(P_{q}^{\text{con}}-\sum_{m\in\mathcal{M}}P_{q,m}^{\text{A}})+\sum_{k\in\mathcal{K}}\left\lVert\mathbf{f}_{k}\right\rVert^{2}$ is the system total consumed power. Constraint set in $\mathcal{R}_{\mathbf{\Theta}}$ in (3b) specifies the feasible region of MF-RISs, i.e., $\alpha_{q,m}\in[0,1]$, $\beta_{q,m}^{k}\in[0,\beta_{\max}^{k}]$, $\theta_{q,m}^{k}\in[0,2\pi)$. (3c) ensures the minimum rate requirement per user as $R^{\min}_{kj}$, while (3d) represents the NOMA power allocation restriction. Constraint (3e) ensures that the total BS transmit power cannot exceed its budget $P_{BS}^{\max}$. (3f) guarantees MF-RIS self-sustainability. Due to non-convexity and non-linearity of problem (3), we propose a DRL-based scheme as detailed in the following section.
@@ III-A Hybrid DRL Algorithm
  • Reward: We design the shared reward as the overall EE in conjunction with its constraints as penalties, given by $r(t)=\frac{\sum_{k\in\mathcal{K}}\sum_{j\in\mathcal{J}_{k}}R_{kj}}{P_{\text{total}}}-\sum_{i=1}^{3}\rho_{i}C_{i}$, where $\rho_{i},\forall i\in\{1,2,3\}$ indicates the weights of each penalty $C_{i}$ corresponding to constraints of (3c), (3e), and (3f), which are defined as $C_{1}=\sum_{k\in\mathcal{K}}\sum_{j\in\mathcal{J}_{k}}(R_{kj}^{\min}-R_{kj})$, $C_{2}=\sum_{k\in\mathcal{K}}\mathbf{\|}\mathbf{f}_{k}\|^{2}-P_{BS}^{\max}$, and $C_{3}=\sum_{q\in\mathcal{Q}}(P_{q}^{\text{con}}-\sum_{m\in\mathcal{M}}P_{q,m}^{\text{A}})$, respectively. Note that the remaining boundary conditions in (3b) and (3d) can be automatically constrained during generating actions.
@@ III-A1 DQN for Discrete Variables
DQN employs a deep neural network, Q-network, to approximate the Q-function $Q_{q}(s_{q},a_{q}|\omega_{q}^{\phi})$ which estimates the expected cumulative reward for each action $a_{q}$ under a given state $s_{q}$. Note that we define $\omega^{\phi}_{q}$ and $\omega^{\phi^{-}}_{q}$ as the model weights of the current network and of the target network of DQN, respectively. Based on estimated Q-values, each agent selects its action using an $\epsilon$-greedy strategy, which balances exploration and exploitation by choosing a random action with probability $\epsilon$ and by selecting the action with the maximum predicted Q-value with probability $1-\epsilon$, i.e., $\epsilon(t)=\epsilon({t-1})-\frac{\epsilon_{\max}-\epsilon_{\min}}{\epsilon_{d}}$, where $\epsilon_{d}$ is the decay parameter. Notations of $\epsilon_{\max}$ and $\epsilon_{\min}$ indicate the maximum and minimum exploration boundaries, respectively. To enhance training stability, DQN incorporates two critical techniques: (i) Experience replay buffer stores historical trajectories with a tuple of $(s_{q},a_{q},r_{q},s^{\prime}_{q})$, allowing the agent to sample mini-batches uniformly and eliminate temporal correlations during learning, where $s^{\prime}_{q}$ indicates the new state; and (ii) Target network, with its model denoted as $Q^{\prime}_{q}(s_{q},a_{q}|\omega_{q}^{\phi^{-}})$ is periodically softly updated to provide stable Q-learning, i.e., $\omega_{q}^{\phi^{-}}\leftarrow\tau_{\phi}\omega_{q}^{\phi}+(1-\tau_{\phi})\omega_{q}^{\phi^{-}}$ where $\tau_{q}^{\phi}$ indicates the importance of target model of DQN. The Q-network is then updated by minimizing the temporal-difference (TD) error, which measures the discrepancy between the predicted Q-value and the target Q-value, given by $\mathcal{L}(\omega_{q}^{\phi})=\mathbb{E}_{(s_{q},a_{q},r,s^{\prime}_{q})}[(y-Q_{q}(s_{q},a_{q}\mid\omega_{q}^{\phi}))^{2}],$ where $y=r(t)+\gamma^{d}\max_{a^{\prime}}Q^{\prime}_{q}(s^{\prime}_{q},a^{\prime}\mid\omega_{q}^{\phi^{-}})$ indicates the TD target value and the discount factor $\gamma^{d}\in[0,1]$ indicates the importance of future rewards. Notation of $Q^{\prime}_{q}(\cdot)$ is the Q-value of the target network. The parameter update via gradient descent is then given by $\omega_{q}^{\phi}\leftarrow\omega_{q}^{\phi}-l^{d}\cdot\nabla_{\omega_{q}^{\phi}}\,\mathcal{L}(\omega_{q}^{\phi})$, where $l^{d}\in[0,1]$ is the learning rate.
@@ III-A2 PPO for Continuous Variables
The remaining continuous parameters are optimized using the PPO algorithm [1]. Particularly, PPO adopts an actor-critic framework respectively consisting of a policy network and of a value network. In the policy network, the neural network outputs the mean and standard deviation of a multivariate Gaussian distribution, from which actions are sampled according to the current state $s_{q}(t)$ and policy $\pi_{\delta_{q}}(a_{q}(t)|s_{q}(t))$. Note that $\delta_{q}$ indicates the policy network parameters. To optimize the policy network, we employ a clipped surrogate objective function expressed as $\mathcal{L}^{\text{clip}}(\delta_{q})={\mathbb{E}}_{t}\big[\min\big(O_{q}(\delta_{q})\hat{A}_{q}(t),\mathrm{clip}(O_{q}(\delta_{q}),1-\Lambda,1+\Lambda)\hat{A}_{q}(t)\big)\big],$ where $\mathbb{E}$$\cdot$] is the expectation over a batch of generated trajectories, and $O_{q}(\delta_{q})=\frac{\pi_{\delta_{q}}(a_{q}(t)|s_{q}(t))}{\pi_{\delta_{q,\text{old}}}(a_{q}(t)|s_{q}(t))}$ is the probability ratio. $\mathrm{clip}(\cdot)$ indicates the clipping function which clips the change between the new and old policies within the range $1-\Lambda,1+\Lambda]$ for avoiding excessive policy updates. Note that $\delta_{q,\text{old}}(\cdot)$ is the old policy parameters. Furthermore, $\hat{A}_{q}(t)$ is the generalized advantage estimation (GAE) quantifying the difference between the observed outcome of each action in a state and the predicted state value $V_{\mu_{q}}(s_{q}(t))$ by the value network, which is $\hat{A}_{q}(t)\!=\!\sum_{i=t}^{T-t}(\gamma^{p}\lambda^{p})^{i-t}\{[(r(i)\!+\!\gamma^{p}V_{\mu_{q}}(s_{q}(i\!+\!1))]\!-\!V_{\mu_{q}}(s_{q}(i))\}$, where $\gamma^{p}$ and $\lambda^{p}$ are importance ratio and GAE hyperparameters, respectively. Notation $T$ means the length of trajectory segment. The policy is optimized iteratively by maximizing the clipped surrogate objective using gradient ascent given by $\delta_{q}\leftarrow{\delta_{q}}-l^{p}_{\text{ac}}\cdot\nabla_{\delta_{q}}\mathcal{L}^{\text{clip}}(\delta_{q})$, where $l^{p}_{\text{ac}}\in[0,1]$ is learning rate for actor in PPO. The associated loss function of the value network is defined as $\mathcal{L}^{V}(\mu_{q})=\mathbb{E}_{t}[\big(V_{\mu_{q}}(s_{q}(t))-\hat{V}^{\text{tar}}_{q}(t)\big)^{2}]$, where $\hat{V}^{\text{tar}}_{q}(t)=V_{\mu_{q}^{-}}\!\big(s_{q}(t)\big)+\hat{A}_{q}(t)$ and $\mu_{q}^{-}$ indicates the previous update of value network. The parameters of value network are updated by the gradient method, i.e., $\mu_{q}\leftarrow{\mu_{q}}-l^{p}_{\text{cr}}\cdot\nabla_{\mu_{q}}\mathcal{L}^{V}(\mu_{q})$, where $l^{p}_{\text{cr}}\in[0,1]$ is learning rate for critic in PPO. Note that coordination is implicitly achieved through shared environment feedback. MF-RIS position adjustments modify the wireless channel state, which is reflected in the observed channel conditions and reward feedbacks from the BS/MF-RIS. As the learning process operates on short decision intervals, the BS/MF-RIS can quickly adapt their strategies based on the updated channels without requiring explicit inter-agent communications.
@@ III-B Parametrized Sharing in PMHRL
Individual agent design in hybrid DRL, PPO and DQN select actions independently based on their own input states. This isolated decision-making process neglects the potential interdependence and interaction between two strategies. To address this, inspired by [21], we propose a parametrized sharing mechanism to enable shared representation between PPO and DQN by exchanging features. Since PPO handling high-dimensional actions performs a more complex task than DQN, information from DQN is essential to PPO. Specifically, PPO agent utilizes the discrete action output from the DQN agent as the input of PPO, i.e., $s_{q}^{\text{con}}(t)=\operatorname{concat}(\mathbf{g}_{q,kj}(t),\operatorname{vec}(a_{q}^{\text{dis}}(t-1)))$, where ${\rm concat}(\cdot)$ indicates the concatenation of two vectors and ${\rm vec}(\cdot)$ vectorizes the discrete action. Note that only MF-RISs require parametrized sharing as they have hybrid actions, thereby improving coordination between the two decision modules.
@@ IV Simulation Results
MF-RIS power consumption parameters [18, 19] $\xi=1.1$, $P_{\text{PIN}}=0.33$ mW, $P_{C}=2.1$ mW, $Z=24$ mW, $\varpi_{1}=150$, $\varpi_{2}=0.014$, $L_{\alpha}=2$, $L_{\beta}=10$, $L_{\theta}=8$
@@
TABLE II: EH Efficiency
$(\varpi_{1},\varpi_{2})$ $(100,0.014)$ $(150,0.014)$ $(200,0.014)$ $(300,0.014)$
Efficiency $0.49$ $0.42$ $0.28$ $0.13$
$(\varpi_{1},\varpi_{2})$ $(150,0.010)$ - $(150,0.018)$ $(150,0.022)$
Efficiency $0.69$ - $0.24$ $0.13$
Figure 2: (a) Convergence. (b) $E$ with different strategies (c) $E$ with different MF-RIS cases. (d) Comparison of NOMA, SDMA, and OMA.
In simulations, we evaluate PMHRL in multi-MF-RISs-assisted downlink NOMA. We consider the BS positioned at $\mathbf{w}_{b}=[0,0,5]$ m, serving $J_{k}=2$ users in $K=2$ directions. Then the users are randomly distributed within a circular area of radius 2 m, centered at $[0,30,0]$, $[0,35,0]$, $[10,40,0]$, and $[10,45,0]$ m, respectively. The MF-RISs are equipped with $M=32$ elements, and the BS has $N=6$ antennas. The remaining parameters related to networks are listed in Table I. As for PMHRL, learning rates of PPO actor/critic networks are $l^{p}_{\text{ac}}=10^{-3}$ and $l^{p}_{\text{cr}}=10^{-4}$, respectively, whereas that of DQN is $l^{d}=10^{-3}$. The discount factor for both modules is $\gamma^{d}=\gamma^{p}=0.99$. The decay and soft update parameters of DQN are set to $\epsilon_{d}=10^{4}$ and $\tau_{\phi}=10^{-2}$, respectively. The experience replay buffer of DQN stores up to $10^{6}$ samples. The mini-batch mechanism is adopted during training, with a batch size of $64$. $\epsilon_{\max}=1$ and $\epsilon_{\min}=0$. Moreover, we set the clipping ratio to $\Lambda=0.2$, the GAE parameter is $\lambda^{p}=0.97$, and trajectory length is $10^{3}$. The weights of each penalty in reward $r(t)$ are $\rho_{1}=10^{-3}$, $\rho_{2}=\rho_{3}=10^{-5}$. Table II shows the EH efficiency with the corresponding constants of circuit sensitivity $\varpi_{1}$ and current leakage $\varpi_{2}$. Higher sensitivity of $\varpi_{1}$ and higher leakage of $\varpi_{2}$ result in a lower EH efficiency value.
@@
Fig. 2 shows that EE escalates with the increasing numbers of antennas thanks to improved beamforming capability, reaching a peak at $N=6$ before declining as the power consumption begins to outweigh beamforming gains. In addition, we compare the fully-optimized case to the cases without either EH ratio $\alpha_{q,m}$, amplification $\beta_{q,m}^{k}$, phase-shift $\theta^{k}_{q,m}$, or deployment $\mathbf{w}_{q}$. Note that ”w/o” indicates random decisions. It is evident that full optimization yields the highest EE. In contrast, omitting the optimization of specific parameters leads to noticeable EE degradation. The configuration without optimizing EH ratio results in the lowest EE, as random S-/H-mode leads to inefficient EH for compensating high power consumption. Moreover, different channel conditions of $\beta_{0}\in\{3,6,12\}$ dB is compared. We can observe that higher $\beta_{0}$ accomplishes higher EE values thanks to better channel conditions with more LoS channel components.
Fig. 2 compares EE of $Q\in\{2,3\}$ MF-RISs under different cases: (1) Optimized case; (2) No EH ($\alpha_{q,m}^{k}=\alpha_{q,m}=1,\forall k\in\mathcal{K}$); (3) No amplification ($\beta_{q,m}^{k}=\beta_{q,m}=1,\forall k\in\mathcal{K}$); (4) Only reflection capability. The results show that increasing MF-RIS elements initially enhances EE. However, deploying more elements induces higher energy consumption, leading to a degraded EE owing to insufficient support from harvested energy. Notably, the case of $Q=3$ MF-RISs outperforms that of $Q=2$ MF-RISs, attributed to the enhanced spatial diversity and EH gain from multi-MF-RISs. Specifically, when signal amplification is disabled, the reduced signal gain leads to a lower EE than that of the fully-optimized case. Moreover, the signal with only reflection cannot be delivered to users beyond the other side, leading to compellingly huge EE reduction. When EH is fully disabled, it cannot provide extra power, resulting in the lowest EE across all cases.
Fig. 2 reveals the EE performance under varying numbers of users per NOMA group and amplification noise levels $\sigma_{s}^{2}\in\{-30,-50,-70,-90\}$ dBm. We compare NOMA to the existing multiple access mechanisms of OMA and spatial division multiple access (SDMA) [9] with or without deployment of MF-RISs. EE decreases with more users per NOMA group due to insufficient power resources and severe inter-user interference. Moreover, NOMA with MF-RIS deployment achieves the highest EE across all numbers of users, benefiting from its superior spectrum utilization and EH capabilities offered by MF-RISs to SDMA and OMA. In contrast, their counterparts without deploying MF-RIS show significantly lower EE performance. Additionally, larger noise power $\sigma_{s}^{2}$ leads to lower EE. Particularly, the amplified noise power of $\sigma_{s}^{2}=-30$ dBm induced by MF-RIS results in the lowest EE and asymptotic performance to that of OMA with MF-RIS.